OpenStudy (shamil98):

Solve the initial-value problem: 2y'' + 5y' + 3y = 0 , y (0) = 3, y'(0) = -4

OpenStudy (shamil98):

The auxillary equation would then be: 2r^2 + 5r + 3 = 0 Using the quadratic formula: r = -3/2 , r = -1 so \[\large y = c_1 e^{-3/2~x} + c_2e^{-x}\] What do i do next?...

OpenStudy (usukidoll):

we can do this either undetermined nevermind ... it's a constant coefficent problem

OpenStudy (usukidoll):

so ok first take the derivative of y that y you just found

OpenStudy (shamil98):

okay. so \[y = -3/2c_1e^{-3/2x} - c_2e^{-x}\] right?

OpenStudy (shamil98):

y'(x)* i should put

OpenStudy (usukidoll):

yeah

OpenStudy (usukidoll):

ok so y(0)=3 is an initial value problem condition so we have it in the form of y(x_o)=yo so the x is equal to 0 and the y is equal to 3 for the ORIGINAL Y

OpenStudy (shamil98):

Oh, so i just plug in!

OpenStudy (usukidoll):

yeah and then find c1 and c2.. afterwards plug in the value of c1 and c2 in the original equation

OpenStudy (usukidoll):

you can check this with laplace if you know how to do it... it would be easy because the laplace transformation of 0 is umm 0 :P

OpenStudy (shamil98):

Not sure i would do it with laplace i'm just skipping around topics with calculus xD

OpenStudy (usukidoll):